# A Categorical Framework for Quantifying Emergent Effects in Network Topology

### v3

@@ 7 Discussion and Conclusion
Emergence, generally conceived as the novel dynamics, patterns or behaviors of a system that are not present in the constituent parts of the system. We adopt a universal and abstract definition of emergence which captures the nonlinearity in system’s effect or observation as we go from components to the whole system. We have established a framework that quantifies the emergence of systems based on their structural information, modeled by quiver representations, and we build a computational measure that quantifies the emergence of systems realized as the mapping between networks where only partial nodes and edges are being preserved, which can be used to model a large variety of systems characterized by the information flow on the network. Other kinds of information flow, for example, those involving coarse-graining, can also be studied by constructing the functors that capture the information flow and compute their derived functors. The emergence is formulated in homological algebra as the ”loss of exactness”, the loss of the algebraic property when a sequence of subsystems modeling interactions is mapped to their partial observations. We showed that our measure correlates with the information-theoretic measure of emergence on random Boolean networks. One important feature of our measure is, as discussed in Section 6, it offers a way to study the contribution of each node to emergence, which is a global property. This could allow us to track the contribution of each node to emergence in a dynamic process, which can help detect the mechanisms of emergence.
@@
From a modeling and computational perspective, our work provides an approach to quantify the emergence of systems as networks, or more generally, quiver representations. It can capture not only the connectivity patterns of the components, but also the structures and dynamics at each node and edge. Within the framework of quiver representations, the mappings between vector spaces need to be linear maps, which renders the framework difficult to capture the non-linearity, like the activation function in the artificial and biological neural networks. But still, quiver representation can be applied to model data and its network structures, so our work can be applied to the phase space of the dynamics, like the spike train data and cellular automata. Also, there has been theoretical development to model neural networks based on quiver representations (see Jeffreys & Lau, 2022; Armenta & Jodoin, 2021). Additionally, the computation of emergence based on networks or quiver representations would often involve counting the number of certain paths in the network, which would require searching algorithms that are time-consuming if the network is large. We call for new approximation schemes and algorithms that evaluate emergence in a cheaper way.
To our knowledge, this is the first work that gives computational measure of emergent effects in a system and its constituent parts, thus opening up exciting new research avenues that lead to an understanding of emergence mechanisms of systems of interest. The computational approach to emergence should appeal to people in a wide range of fields where there is a notion for emergence. It should be mentioned here that the term ”generativity” can almost be used interchangeably with emergence. We choose to use emergence since it is a term widely used across fields while ”generativity” is mainly discussed in artificial intelligence.

### v4

@@ 7 Discussion and Conclusion
Emergence, generally conceived as the novel dynamics, patterns or behaviors of a system that are not present in the constituent parts of the system. We adopt a universal and abstract definition of emergence which captures the structural nonlinearity in system’s effect or observation as we go from components to the whole system. We have established a framework that quantifies the emergence of systems based on their structural information, modeled by quiver representations, and we build a computational measure that quantifies the emergence of systems realized as the mapping between networks where only partial nodes and edges are being preserved, which can be used to model a large variety of systems characterized by the information flow on the network. Other kinds of information flow, for example, those involving coarse-graining, can also be studied by constructing the functors that capture the information flow and compute their derived functors. The emergence is formulated in homological algebra as the ”loss of exactness”, the loss of the algebraic property when a sequence of subsystems modeling interactions is mapped to their partial observations. We showed that our measure correlates with the information-theoretic measure of emergence on random Boolean networks. One important feature of our measure is, as discussed in Section 6, it offers a way to study the contribution of each node to emergence, which is a global property. This could allow us to track the contribution of each node to emergence in a dynamic process, which can help detect the mechanisms of emergence.
@@
From a modeling and computational perspective, our work provides an approach to quantify the emergence of systems as networks, or more generally, quiver representations. It can capture not only the connectivity patterns of the components, but also the structures and dynamics at each node and edge. Within the framework of quiver representations, the mappings between vector spaces need to be linear maps, which renders the framework difficult to capture the non-linearity, like the activation function in artificial and biological neural networks. But still, quiver representation can be applied to model data and its network structures, so our work can be applied to the phase space of the dynamics, like the spike train data and cellular automata. Also, there has been theoretical development to model neural networks based on quiver representations (see Jeffreys & Lau, 2022; Armenta & Jodoin, 2021). Additionally, the computation of emergence based on networks or quiver representations would often involve counting the number of certain paths in the network, which would require searching algorithms that are time-consuming if the network is large. We call for new approximation schemes and algorithms that evaluate emergence in more efficient ways.
To our knowledge, this is the first work that gives a computational measure of emergent effects in a system and its constituent parts, thus opening up exciting new research avenues that lead to an understanding of the emergence mechanisms of systems of interest. The computational approach to emergence should appeal to people in a wide range of fields where there is a notion for emergence. It should be mentioned here that the term ”generativity” can almost be used interchangeably with emergence. We choose to use emergence since it is a term widely used across fields while ”generativity” is mostly discussed in artificial intelligence.